A 50-g piece of metal at 95°C is dropped into 250g of water at 17°C and warms it to 19.4°C. What is the specific heat of the metal?


1
Expert's answer
2022-04-01T16:32:38-0400

Given:

"t_1=95^\circ\rm C"

"t_2=17^\circ\rm C"

"t_3=19.4^\circ\rm C"

"m_1=50\:\rm g"

"m_2=250\:\rm g"

The heat balance gives


"Q_1=Q_2\\
cm_1(t_1-t_3)=c_wm_2(t_3-t_2)"

Hence, the specific heat of the metal

"c=c_w\frac{m_2(t_3-t_2)}{m_1(t_1-t_3)}"

"c=\rm4.186\: J/g*\frac{250(19.4-17)}{50(95-19.4)}=0.664\: J/g"